OpenStudy (anonymous):

simplify -7 1/8 - (-9 1/2) 1. -2 3/8 2. -16 5/8 3. 16 5/8 4. 2 3/8

OpenStudy (anonymous):

4

OpenStudy (anonymous):

how

OpenStudy (anonymous):

i dont want to write out or draw it sorry

OpenStudy (anonymous):

but its right

OpenStudy (michele_laino):

hint: we can write: \[\Large \begin{gathered} - 7\frac{1}{8} = \frac{{ - 7 \times 8 - 1}}{8} = - \frac{{57}}{8} \hfill \\ \hfill \\ - 9\frac{1}{2} = \frac{{ - 9 \times 2 - 1}}{2} = - \frac{{19}}{2} \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

so, we have: \[ \Large - 7\frac{1}{8} - \left( { - 9\frac{1}{2}} \right) = - \frac{{57}}{8} - \left( { - \frac{{19}}{2}} \right) = - \frac{{57}}{8} + \frac{{19}}{2} = ...\]

OpenStudy (michele_laino):

\[\large - 7\frac{1}{8} - \left( { - 9\frac{1}{2}} \right) = - \frac{{57}}{8} - \left( { - \frac{{19}}{2}} \right) = - \frac{{57}}{8} + \frac{{19}}{2} = ...\]